OpenStudy (anonymous):

what is x when 9x+4x+37=180

OpenStudy (anonymous):

Combine terms first. You can add the 9x and 4x together because they are "like terms." You can then subtract 37 from both sides of the equation. So here is what we have: 13x = 143 Then divide both sides of the equation by 12 and you get x=143/13 = 11 x=11

OpenStudy (anonymous):

13x=143 x=143/13 x=11

OpenStudy (anonymous):

Oops meant divide both sides of the equation by 13. Typo.

OpenStudy (anonymous):

13x+37=180 13x=143 x=143/13 x=11